[figure: 9.7 cm]

TUNNELING CURRENT IN NORMAL METAL- ODD PARITY

SUPERCONDUCTOR JUNCTION

Leon Kowalewski, Ryszard J. Wojciechowski and Piotr Wojtuś A. Mickiewicz University, Institute of Physics ul. Umultowska 85, 61-614 Poznań, Poland

The properties of tunneling conductance spectra of a normal metal- insulator- spin triplet superconductor junction are investigated. We take into account a higher (p-wave and fwave)relative angular momentum contributions to the superconducting order parameter and investigate the quasiparticle tunneling for both the unitary and non-unitary pairing states, which are consistent with the experimental data for Sr_{2}RuO_{4}. The zero-bias conductance peaks due to the quasiparticle bound states at the interface boundary and the gap-like structure of the conductance depending on the orientation of the junction are obtained and the role of the Andreev reflection in indentifying the pairing symmetry in the odd-parity superconductors is also discussed.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Ryszard J. Wojciechowski

Address for correspondence : A. Mickiewicz University, Institute of Physics ul. Umultowska 85, 61-614 Poznań, Poland

Email address : wojrysz@main.amu.edu.pl